OpenStudy (anonymous):

What is the domain of function √x+4/3-x

geerky42 (geerky42):

\[\sqrt{\dfrac{x+4}{3-x}}\]?

OpenStudy (gorv):

denominator should not be equal to zero ...that is \[x \neq 3\]

OpenStudy (gorv):

\[\frac{ x+4 }{ 3-x }\ge0\]

OpenStudy (gorv):

solve it